In need of extra​ cash, Troy and Lilly decide to withdraw ​$9,400 from their traditional IRA. They are both 40 years old. They are in a 22​% marginal tax bracket. What will be the tax consequences of this​ withdrawal? ​(Note​: Assume a penalty of​ 10% of the withdrawn​ funds.) The tax consequences of this withdrawal will be ​$___
Answer #1

Early withdrawal from Traditional IRA will be treated as income ( not capital gain ) and charged to as such. A penalty of 10% and applicable marginal tax rate will be levied to compute tax on such withdrawal. However such withdrawal will be exempt if it is withdrawn for certain reasons however need of extra money is not one of them. Hence $9,400 will be charged by 10% penalty and 22% tax.

The tax consequences of this withdrawal will be ​$3,008

Early withdrawal penalty = 9,400 * 10 % = 940

Taxes = 9,400 * 22% = 2,068

Total= $3,008